OpenStudy (anonymous):

Medal! Use the distributive property to rewrite the expression and solve. 8 63 = 8 ? + 8 3 = ?

jimthompson5910 (jim_thompson5910):

hint: 63 = 60+3

jimthompson5910 (jim_thompson5910):

the distributive property is a*(b+c) = a*b + a*c

OpenStudy (anonymous):

Is 8*63=8*60+8*3=504 right?

jimthompson5910 (jim_thompson5910):

you nailed all 3 parts

jimthompson5910 (jim_thompson5910):

nice work

jimthompson5910 (jim_thompson5910):

although I would add in one more step the part about 8*(60+3)

jimthompson5910 (jim_thompson5910):

so 8*63=8*(60+3)=8*60+8*3=504

OpenStudy (anonymous):

Thank you for the help!

jimthompson5910 (jim_thompson5910):

but to keep to the format `8 63 = 8 ? + 8 3 = ?` your answer is perfect

jimthompson5910 (jim_thompson5910):

np